OpenStudy (ac3):

A diode laser emits at wavelength of 744nm. All of its output energy is absorbed in a detector that measures a total energy of .63J over a period of 47s. How many photons per second are being emitted by the laser. I got 5.0*10^16 photons/second. Back of the book says 5.0*10^13photons/second.

OpenStudy (ac3):

alright one sec.

OpenStudy (ac3):

E=hv... h being planck's constant. v being frequency.

OpenStudy (ac3):

\[\lambda v = c\]

OpenStudy (ac3):

so \[v=c/\lambda \]

OpenStudy (ac3):

\[E=h(c/\lambda )\]

OpenStudy (ac3):

\[E=6.626^{-34}Js*3*10^{8}m/s*\frac{ 1 }{ 744*10^{-9}m }\]

OpenStudy (ac3):

\[2.67*10^{-19}\]

OpenStudy (ac3):

thats what tha equals.

OpenStudy (ac3):

and that's in J/photon

OpenStudy (ac3):

\[\frac{ .63J }{ 47s}*\frac{ photon }{ 2.67*10^{-19} J}\]

OpenStudy (ac3):

\[=5.0*10^{16}\]

OpenStudy (ac3):

photons/second

OpenStudy (ac3):

@Rushwr that's how I went about it i'm assuming that's the process not sure if i plugged in to the calculator properly.

OpenStudy (ac3):

@IrishBoy123 you good with chemistry?

OpenStudy (ac3):

@freckles you good with chem?

OpenStudy (freckles):

no but I can try to look at your problem

OpenStudy (ac3):

try and just plug in the numbers into a calculator and see if you get the correct answer

OpenStudy (freckles):

I have done that and it doesn't calculate to the back of the book \[E=6.626^{-34}Js*3*10^{8}m/s*\frac{ 1 }{ 744*10^{-9}m }\] where does the 6.626^(-34) come from and the 3*10^8 come from ?

OpenStudy (ac3):

6.636*10^-34 is planck's constant.

OpenStudy (ac3):

which is denoted using h.

OpenStudy (freckles):

oh times 10^(-34)

OpenStudy (ac3):

and c is the speed of light.

OpenStudy (ac3):

which is the 3*10^8m/s

OpenStudy (freckles):

I do have one complaint

OpenStudy (ac3):

when you get that first part the units are in joules/photon and then you continue using dimensional analysis to get photon/second

OpenStudy (ac3):

what is it?

OpenStudy (freckles):

\[7.44 \times 10^{-7} m =744 nm\] right?

OpenStudy (ac3):

yes

OpenStudy (freckles):

you have -9

OpenStudy (freckles):

not -7

OpenStudy (ac3):

i account for it by saying 744 though

OpenStudy (ac3):

i kept the decimal where it was given and put 10^-9m

OpenStudy (ac3):

which is what a nanometer equals

OpenStudy (freckles):

oh ok

OpenStudy (freckles):

ok give me a few i'm thinking

OpenStudy (ac3):